fred is 3 times older than his sister Mary, four years ago he was 5 times as old as his sister. find their present ages
Answer by macston(5194)

.
M=Mary's age; F=Fred's age=3M
.
F-4=5(M-4)
3M-4=5M-20
16=2M
M=8 ANSWER 1: Mary is 8 years old.
F=3M=3(8)=24 ANSWER 2: Fred is 24 years old.
.
CHECK:
F-4=5(M-4)
24-4=5(8-4)
20=3(4)
20=20